OpenStudy (anonymous):

A 125-ft diagonal brace on a bridge connects a support of the center of the bridge to a side support on the bridge.The horizontal distance that it spans is 25 ft longer than the height that it reaches on the side of the bridge. Find the horizontal and vertical distances spanned by this brace.

OpenStudy (saifoo.khan):

\[x^2 + (x+25)^2 = 125^2\]

OpenStudy (kira_yamato):

Let the vertical distance be x ft. Then the horizontal distance is (x + 25) ft. Therefore, by Pythagorean theorem, x^2 + (x + 25)^2 = 125^2 Solve for x, and then solve for (x + 25)

OpenStudy (anonymous):

thank you I think I can figure this out now.